i want to make a solution of sodium chloride of 0.2 moll but the volume has to 50 cm cubed. How would i make the calculations to find out how much of everything is needed?
1
Expert's answer
2013-01-16T11:58:01-0500
0.2 mol/ L means that solution with volumeof 1 L dissolved 0.2 mol of NaCl

if you have 0.2 mol in 1000 cm^3
you'll have x mol in 50 cm^3

x = 0.01 mol

n = m/Mw

m = n*Mw, Mw of NaCl is 58.44

m = 0.01*58.44 =0.5844 g

So you need to mix 0.5844 g of NaCl and 50 cm^3 (50 ml) of
water.